Q1) Write an algebraic expression. Then write an equation for the variable indicated for the last sentence.
Let b be Bonnie's age now.
a) What was her age 6 years ago?
b) If Clyde is 1 year younger than Bonnie, represent his age now.
c) What was Clyde's age 6 years ago?
d) If 6 years ago the sum of their ages was 35, write an equation to represent
this.

Answer by ikleyn(52777)

.
Q1) Write an algebraic expression. Then write an equation for the variable indicated for the last sentence.
Let b be Bonnie's age now. // and c be the Clyde's age now.
 a) What was her age 6 years ago?                                     b-6.
 b) If Clyde is 1 year younger than Bonnie, represent his age now.    c = b-1.
 c) What was Clyde's age 6 years ago?                                 c-6
 d) If 6 years ago the sum of their ages was 35, write an equation  
    to represent this.                                                (b-6) + (c-6) = 35,   or, substituting c = b-1 from above,
                                                                      (b-6) + (b-7) = 35,   which is  2b-13 = 35.